# myequation

### v1

@@ Abstract
This work introduces a lightweight simulation framework for evaluating asynchronous entanglement distribution protocols under realistic error models. We focus on two contemporary protocols: sequential, where entanglement is established one node at a time, and parallel, where all nodes attempt to generate entanglement simultaneously. We evaluate the performance of each protocol using two key metrics: the fidelity of distributed entangled states, and the hashing rate, a measure of entanglement efficiency. These metrics are compared between both protocols across a range of network sizes and noise parameters. We demonstrate that the parallel protocol consistently outperforms the sequential, particularly in the hashing rate metric due to reduced runtime, suggesting that parallel protocols are a strong candidate for a realizable quantum Internet. Our framework offers an accessible and scalable tool for evaluating entanglement distribution strategies, by reducing the simulation of complex quantum processes to simple memory time calculations.
@@ 2 Entanglement Distribution and Noise in Quantum Networks
The process of Bell state distribution begins by selecting two end nodes, one of which serves as the sender node (which initiates the communication), and the other as the receiver node (which receives the signal). Between these two end nodes exists a collection of intermediate repeater nodes, each of which serves to mitigate optical loss occurring throughout the protocol. We first consider the 222-qubit Bell state ${|{\phi^{+}}\rangle}$, defined
@@
[…]
@@ 2.1 An Error Model for Entanglement-Based Quantum Networks
$\eta_{ch}(L)\equiv e^{\frac{L}{L_{att}}},$ (2)
where $L$ denotes the length the photon must travel between nodes, and $L_{att}$ is the attenuation length. For optical fibers operating at Telecom wavelengths, the primary focus of this work, the value of $L_{att}$ is typically around 222222 km.
@@
Successful entanglement swapping depends on the ability to perform high-fidelity Bell state measurement (BSM) on transmitted quantum states. A Bell state measurement consists of a 222-qubit measurement, followed by the projection of their combined state onto one of four maximally entangled Bell states, specifically
$\begin{split}{|{\phi^{\pm}}\rangle}&\equiv\frac{1}{2}\left({|{00}\rangle}\pm{|{11}\rangle}\right),\\ {|{\psi^{\pm}}\rangle}&\equiv\frac{1}{2}\left({|{01}\rangle}\pm{|{10}\rangle}\right).\\ \end{split}$ (5)
When performing entanglement swapping on noisy hardware, perfectly distinguishing each Bell state in Eq. (5) and performing the necessary projective measurement is rarely achieved without error. Imperfect Bell state measurement occurs due to a variety of factors, including photon detector limitations, decoherence, and the indistinguishability of states in certain optical realizations. In this work, we model imperfect Bell state measurement by applying a 222-qubit depolarizing channel $\mathcal{E}_{w}^{(i,j)}$ to the qubits $i$ and $j$ being measured, followed by an ideal Bell state measurement. We define the depolarizing channel $\mathcal{E}_{w}^{(i,j)}$ as
@@ 2.2 A Closed Form Expression for Fidelity in an $n$-Qubit Network
In this section, we derive a closed expression for the fidelity of an entangled state distributed by an $n$-qubit quantum network using the error model presented in the previous section. Throughout this work, we use $n$ to denote the number of qubits and $N=2n-1$ to represent the number of network nodes. We first demonstrate that dephasing error depends only on the total time spent in memory across all qubits, and grows exponentially as the ratio of this cumulative time over the memory dephasing time. We then combine the effect of imperfect Bell state measurement error to arrive at a general expression for the fidelity of a Bell state prepared using entanglement swapping in an $n$-qubit quantum network.
To demonstrate that the dephasing error depends only on the total time spent in memory, we firstly consider a quantum network of four qubits. The network consists of two Bell pairs $({|{\phi^{+}}\rangle}\bra{\phi^{+}})_{AB}$ and $({|{\phi^{+}}\rangle}\bra{\phi^{+}})_{CD}$, shared between qubits $A$ and $B$, and qubits $C$ and $D$ respectively. We allow each qubit to remain in memory for some fixed amount of time. By performing a Bell state measurement on qubits $B$ and $C$, a new Bell pair, $({|{\phi^{+}}\rangle}\bra{\phi^{+}})_{AD}$, can be generated. The effect of dephasing on the fidelity $F_{AD}$ of the distributed Bell pair is derived in Appendix A.1, and is given by
@@
Extending this result further, we now consider a 666-qubit network comprised of three Bell pairs $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{AB}$, $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{CD}$, and now $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{EF}$ as well. We first perform an entanglement swapping event, as before, generating $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{AD}$. We then implement entanglement swapping between $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{AD}$ and $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{EF}$, performing a Bell state measurement on qubits $D$ and $E$, resulting in the state $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{AF}$. The fidelity $F_{AF}$ of the distributed Bell state, $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{AF}$, is derived in Appendix A.2, and has the form
@@
Since, for a repeater quantum network, expanding to a higher number of nodes involves iteratively appending Bell pairs and performing subsequent entanglement swapping operations, the fidelity expressions derived in Eqs. (10) and (13) apply generically for all $n$. Otherwise stated, given a repeater quantum network of $n$ qubits $\{a_{1},...,a_{n}\}$, the Fidelity of a Bell state $\left({|{\phi^{+}}\rangle}\bra{\phi^{+}}\right)_{a_{1},a_{n}}$ generated between $a_{1}$ and $a_{n}$ via entanglement swapping events is
@@
where $m$ denotes the number of Bell state measurements needed to perform all entanglement swaps. In linear networks $m=N-2$, i.e. entanglement swapping must be performed on all quantum repeaters in the $N$-node network. Eq. (17) gives a general expression for a Bell state prepared using $m$ entanglement swapping events in an $n$-qubit repeater network.
In this section, we derived a general closed expression for the fidelity of a Bell state prepared via entanglement swapping in an $n$-qubit quantum network which incorporates the errors of fiber loss, memory dephasing and imperfect Bell state measurements. Firstly, we demonstrated that the dephasing error depends only on the total time all qubits spend in memory, rather than on the distribution of individual quantum memory times. This was demonstrated through an expression for the fidelity of the final distributed Bell state with the only variables being the total time in memory, $T$, and the memory dephasing time $T_{dp}$. The total time spent in memory during the protocol is affected by fiber loss and can be obtained through simple, classical simulation. Our final fidelity expression, Eq. (17), accounts for depolarization from imperfect Bell state measurements as well as dephasing error from quantum memories.
@@ 2.3 Performance Evaluation Metrics
where $dn/dt$ represents the rate at which imperfect Bell pairs are generated by the quantum network. Together, Eqs. (18) and (20) determine the overall success and efficiency of entanglement distribution and distillation in an $n$-qubit repeater network.
@@
In this section, we introduced success metrics for evaluating the performance of the entanglement distribution simulations for quantum repeater networks. We used fidelity as a closeness measure between the network output state and the desired Bell state, and hashing rate to quantify the efficiency of distilling Bell-type entanglement from imperfect Bell pairs produced by the quantum network. We reviewed the respective constructions of parallel and sequential entanglement distribution protocols, which we now simulate in the following section.
3 Efficient Simulation of Asynchronous Entanglement Distribution
In this section we simulate entanglement distribution using the parallel and sequential protocols, incorporating the noise model derived in Section 2. Leveraging insight from our analytical result in Eq. (17), we observe that the fidelity of the final Bell state depends on several constant parameters specific to a given network setup, namely the number of Bell state measurements $m$, the BSM ideality parameter $\lambda_{\text{BSM}}$, and the memory dephasing time $T_{\text{dp}}$. Crucially, among these factors, only the cumulative time that qubits spend in memory $T$, varies across different runs of the protocol in the same network. As a result, we can significantly simplify our simulations by computing just two key quantities: the total quantum memory time $T$ to calculate fidelity, and the overall protocol duration $T_{\text{clock}}$ to calculate hashing rate. We present and analyze the results of our simulation, and examine the role of imperfect Bell state measurement in each case. Finally, we compare the strengths and limitations of each protocol, and highlight their respective advantages and drawbacks for practical entanglement distribution in near-term quantum networks.
The simulations in this work are conducted under a base set of assumptions. We outline the key assumptions below that define our simulation framework.
@@ 3 Efficient Simulation of Asynchronous Entanglement Distribution
We track the entanglement links between nodes in the quantum network state using a Boolean array of length $N-1$, where $N$ is the total number of nodes. Each array element corresponds to a link between adjacent nodes, and carries a value “True” if entanglement has been successfully established between the links adjacent nodes. For our simulation, time is normalized to $L/c$, where $L$ is the internode distance and $c$ is the speed of light in optical fiber. We first consider only the error due to dephasing in quantum memory. Since dephasing and BSM errors are independent this simplification is justified. In Section 3.2, we extend our analysis to include the impact of BSM imperfections.
Given that the final Bell state fidelity depends solely on the cumulative time qubits spend in memory, we can characterize the performance of both the sequential and parallel distribution protocols by analyzing the total quantum memory time $T$ and the overall protocol duration $T_{\text{clock}}$. The latter is required specifically for calculating the hashing rate. We now detail how both $T$ and $T_{\text{clock}}$ are computationally obtained in the simulations, and provide an algorithmic implementation for each protocol. All code used in these simulations is publicly available online [33].
@@
  • $T_{\text{clock-step}}=2$: Entanglement generation events take 222 time steps to complete.
  • $T_{\text{success}}=3$: Cumulative memory time acquired by successfully generating an entanglement link. When generating entanglement between $A$ and $B$, Node $A$ adds 222 time steps to memory, Node $B$ adds 111 time step (Section 2).
@@
[…]
3:links $=$ [False] $\times$ $(N-1)$
@@
[…]
7:     link $=$ True
8:     if link is not first then $T\mathrel{+}=T_{\text{existing}}$
9:     $T_{\text{clock}}\mathrel{+}=T_{\text{clock-step}}$
10:return $(T,\,T_{\text{clock}})$
The sequential procedure iteratively attempts to establish entanglement between each pair of adjacent nodes along a linear chain of repeaters. At each step, entanglement is attempted until successful, the probability of successful entanglement generation is governed by a loss parameter $p_{\text{loss}}$. If successful, a memory time of $T_{\text{success}}$ is added to the cumulative memory time $T$. For all links except the first, an additional penalty $T_{\text{existing}}$ is added to account for decoherence of the two qubits held in memory as the entanglement attempt is taking place. The clock is incremented by $T_{\text{clock-step}}$ per attempt. The algorithm terminates once all $N-1$ links are successfully established, returning the total memory time $T$ and the total elapsed time $T_{\text{clock}}$.
@@
[…]
3:links $=[\text{False}]\times(N-1)$
4:while not AllSwapped(linksdo
@@
3:  for each index $i$ in $0\,\dots\,N-2$ do
[…]
7:      $T_{\text{round}}+=T_{\text{success}}$
@@
In the parallel scheme, all links between adjacent nodes in a linear chain attempt entanglement generation simultaneously during each round. The probability of successful entanglement generation is again governed by the loss parameter $p_{\text{loss}}$. In successful entanglement generation events, the link value is updated to “True”, and a time of $T_{\text{success}}$ is added to the cumulative memory time $T$. An additional memory time of 444 units is added to $T_{\text{round}}$ for each run of established links to account for two qubits held in memory per round of entanglement generation attempts. The total clock time $T_{\text{clock}}$ increases by a fixed increment $T_{\text{clock-step}}$ per round. The protocol repeats until all $N-1$ links have been successfully established, returning both the total memory time $T$ and the total elapsed time $T_{\text{clock}}$.
@@ 3.1 Comparison of Sequential and Parallel Protocols
We now evaluate and compare the respective performances of sequential and parallel asynchronous protocols in a linear quantum network of total length 505050 km. Throughout each simulation, we vary both the number of nodes and the memory dephasing time. As discussed in Section 2.3, key metrics for success are the final fidelity of the distributed Bell state and the hashing rate. The results of our simulations are compiled in Figures 1 and 2.
@@
Figure 1: Subfigures (a) and (b) show the fidelity of a distributed Bell pair in the sequential and parallel protocols, respectively. Subfigures (c) and (d) depict corresponding hashing rates, i.e. the rate of high-fidelity Bell pair distillation. While fidelity is comparable across both protocols, the parallel approach demonstrates a strong advantage in hashing rate.
@@
In this section, we presented simulations comparing the performance of sequential and parallel asynchronous entanglement distribution protocols. We evaluated the output fidelity of distributed Bell pairs, with increasing network size and increasing quantum memory dephasing time, as well as the hashing rates achieved by each protocol. While both approaches produced similar fidelities, the parallel protocol demonstrated a clear advantage in hashing rate due to its overall shorter runtime. In the next section we consider the effect of imperfect Bell state measurement on entanglement distribution, and assess its impact on fidelity and hashing rate.

### v2

# A Compact Framework for Analyzing Asynchronous Entanglement Distribution in Quantum Networks

@@ Abstract
This work introduces a compact framework for analyzing asynchronous entanglement distribution protocols under realistic error models. We focus on two contemporary protocols: sequential, where entanglement is established one node at a time, and parallel, where all nodes attempt to generate entanglement simultaneously. We derive an analytical expression for the fidelity of distributed entangled states, showing that the fidelity depends only on the total time all qubits spend in memory, rather than the individual memory times for each qubit. This result distills the complex dynamics of entanglement distribution into a compact accessible form, providing an scalable tool for evaluating protocol efficiency. Using this lightweight framework, we analyze the performance of parallel and sequential protocols, demonstrating that parallel distribution consistently outperforms sequential and highlighting the potential of parallel protocols for practical quantum network implementations.
@@ 2 Entanglement Distribution and Noise in Quantum Networks
The process of Bell state distribution begins by selecting two end nodes, one of which serves as the sender node (which initiates the communication), and the other as the receiver node (which receives the signal). Between these two end nodes exists a collection of intermediate repeater nodes, each of which serves to mitigate optical loss occurring throughout the protocol. We first consider the $2$-qubit Bell state $|{\phi^{+}}\rangle$}, defined
@@
Entanglement swapping involves two pairs of entangled photons $(A,B)$ and $(C,D)$. A Bell state measurement is performed on photons $(B,C)$, resulting in the two shorter-distance entanglements to be effectively swapped into one longer distance entanglement between photons $(A,D)$. Once entanglement is successfully established on both sides of the node, i.e. the node has both successfully received an entangled photon and sent an entangled photon to its neighboring node, an entanglement swapping operation is performed, establishing entanglement between its neighboring nodes.
@@ 2.1 An Error Model for Entanglement-Based Quantum Networks
$\eta_{ch}(L)\equiv e^{\frac{-L}{L_{att}}},$ (2)
where $L$ denotes the length the photon must travel between nodes, and $L_{att}$ is the attenuation length. For optical fibers operating at Telecom wavelengths, the primary focus of this work, the value of $L_{att}$ is typically around $22$ km.
@@
Successful entanglement swapping depends on the ability to perform high-fidelity Bell state measurement (BSM) on transmitted quantum states. A Bell state measurement consists of a $2$-qubit measurement, followed by the projection of their combined state onto one of four maximally entangled Bell states, specifically
[…]
When performing entanglement swapping on noisy hardware, perfectly distinguishing each Bell state in Eq. (5) and performing the necessary projective measurement is rarely achieved without error. Imperfect Bell state measurement occurs due to a variety of factors, including photon detector limitations, decoherence, and the indistinguishability of states in certain optical realizations. In this work, we model imperfect Bell state measurement by applying a $2$-qubit depolarizing channel $\mathcal{E}_{w}^{(i,j)}$ to the qubits $i$ and $j$ being measured, followed by an ideal Bell state measurement. We define the depolarizing channel $\mathcal{E}_{w}^{(i,j)}$ as
@@ 2.2 A Closed Form Expression for Fidelity in an $n$-Qubit Network
In this section, we derive a closed expression for the fidelity of an entangled state distributed by an $n$-qubit quantum network using the error model presented in the previous section. Throughout this work, we use $N$ to denote the number of network nodes and n=2(N1)n=2(N-1) to represent the number of qubits. We first demonstrate that dephasing error depends only on the total time spent in memory across all qubits, and grows exponentially as the ratio of this cumulative time over the memory dephasing time. We then combine the effect of imperfect Bell state measurement error to arrive at a general expression for the fidelity of a Bell state prepared using entanglement swapping in an $n$-qubit quantum network.
[…]
@@
[…]
@@
[…]
@@
where $m$ denotes the number of Bell state measurements needed to perform all entanglement swaps. In linear networks m=N2m=N-2, i.e. entanglement swapping must be performed on all quantum repeaters in the $N$-node network.
Eq. (17) provides a general expression for Bell state fidelity, prepared using $m$ entanglement swapping events in an $n$-qubit repeater network. The dependence of this fidelity expression on the total memory time $T$, as opposed to individual qubit memory times, significantly simplifies the computation. Instead of relying on memory timing histories to compute fidelity, as in previous work [27], Eq. (17) provides a compact analytical formulation that enables efficient analysis at substantially larger qubit scales, as we demonstrate in Section 3.1.
In this section, we derived a general closed expression for the fidelity of a Bell state prepared via entanglement swapping in an $n$-qubit quantum network which incorporates the errors of fiber loss, memory dephasing and imperfect Bell state measurements. Firstly, we demonstrated that the dephasing error depends only on the total time all qubits spend in memory, rather than on the distribution of individual quantum memory times. This was demonstrated through an expression for the fidelity of the final distributed Bell state with the only variables being the total time in memory, $T$, and the memory dephasing time $T_{dp}$. The total time spent in memory during the protocol is affected by fiber loss and can be obtained through simple classical simulation. Our final fidelity expression, Eq. (17), accounts for depolarization from imperfect Bell state measurements as well as dephasing error from quantum memories.
@@ 2.3 Performance Evaluation Metrics
where $\frac{dn}{dt}$ is the rate at which imperfect Bell pairs are generated across the quantum network. It is defined as
$\frac{dn}{dt}=\frac{1}{T_{clock}}$ (21)
where $T_{clock}$ represents the total time taken to distribute a Bell pair across the network. Together, Eqs. (18) and (20) determine the overall success and efficiency of entanglement distribution and distillation in an $n$-qubit repeater network.
@@
In this section, we introduced success metrics for evaluating the performance of the entanglement distribution protocols for quantum repeater networks. We used fidelity as a closeness measure between the network output state and the desired Bell state, and hashing rate to quantify the efficiency of distilling Bell-type entanglement from imperfect Bell pairs produced by the quantum network. We reviewed the respective constructions of parallel and sequential entanglement distribution protocols, which we now simulate in the following section.
3 Efficient Analysis of Asynchronous Entanglement Distribution
In this section we simulate entanglement distribution using the parallel and sequential protocols, incorporating the noise model derived in Section 2. Leveraging insight from our analytical result in Eq. (17), we observe that the fidelity of the final Bell state depends on several constant parameters specific to a given network setup, namely the number of Bell state measurements $m$, the BSM ideality parameter $\lambda_{\text{BSM}}$, and the memory dephasing time $T_{\text{dp}}$. Crucially, among these factors, only the cumulative time that qubits spend in memory $T$, varies across different runs of the protocol in the same network. As a result, we can significantly simplify our analysis by computing just two key quantities: the total quantum memory time $T$ to calculate fidelity, and the overall protocol duration $T_{\text{clock}}$ to calculate hashing rate. We present and analyze the results of our analysis, and examine the role of imperfect Bell state measurement in each case. Finally, we compare the strengths and limitations of each protocol, and highlight their respective advantages and drawbacks for practical entanglement distribution in near-term quantum networks.
The analysis in this work are conducted under a base set of assumptions. We outline the key assumptions below that define our analytical framework.
@@ 3 Efficient Analysis of Asynchronous Entanglement Distribution
We track the entanglement links between nodes in the quantum network state using a Boolean array of length N1N-1, where $N$ is the total number of nodes. Each array element corresponds to a link between adjacent nodes, and carries a value “True” if entanglement has been successfully established between the links adjacent nodes. For our analysis, time is normalized to $L/c$, where $L$ is the internode distance and $c$ is the speed of light in optical fiber. We first consider only the error due to dephasing in quantum memory. Since dephasing and BSM errors are independent this simplification is justified. In Section 3.2, we extend our analysis to include the impact of BSM imperfections.
Given that the final Bell state fidelity depends solely on the cumulative time qubits spend in memory, we can characterize the performance of both the sequential and parallel distribution protocols by analyzing the total quantum memory time $T$ and the overall protocol duration $T_{\text{clock}}$. The latter is required specifically for calculating the hashing rate. We now detail how both $T$ and $T_{\text{clock}}$ are computationally obtained in the analysis, and provide an algorithmic implementation for each protocol. All code used in this work is publicly available online [33].
@@
  • $T_{\text{clock-step}}=2$: Entanglement generation events take $2$ time steps to complete.
  • $T_{\text{success}}=3$: Cumulative memory time acquired by successfully generating an entanglement link. When generating entanglement between $A$ and $B$, Node $A$ adds $2$ time steps to memory, Node $B$ adds $1$ time step (Section 2).
@@
3:links == [False] $\times$ (N1)(N-1)
@@
6:   if RandomSuccess($p_{\text{loss}}$) then
7:     $T\mathrel{+}=T_{\text{success}}$
8:     link == True
9:     if link is not first then
10:      $T\mathrel{+}=T_{\text{existing}}$
11:   $T_{\text{clock}}\mathrel{+}=T_{\text{clock-step}}$
12:return $(T,\,T_{\text{clock}})$
The sequential procedure iteratively attempts to establish entanglement between each pair of adjacent nodes along a linear chain of repeaters. At each step, entanglement is attempted until successful, the probability of successful entanglement generation is governed by a loss parameter $p_{\text{loss}}$. If successful, a memory time of $T_{\text{success}}$ is added to the cumulative memory time $T$. For all links except the first, an additional penalty $T_{\text{existing}}$ is added to account for decoherence of the two qubits held in memory as the entanglement attempt is taking place. The clock is incremented by $T_{\text{clock-step}}$ per attempt. The algorithm terminates once all N1N-1 links are successfully established, returning the total memory time $T$ and the total elapsed time $T_{\text{clock}}$.
@@
3:links $\text{False}]\times(N-1)$
4:while not AllSwapped(links) do
@@
3:  for link in links do
4:   if link = False then
5:     if RandomSuccess($p_{\text{loss}}$) then
6:      link = True
7:      $T_{\text{round}}\mathrel{+}=T_{\text{success}}$
@@
In the parallel scheme, all links between adjacent nodes in a linear chain attempt entanglement generation simultaneously during each round. The probability of successful entanglement generation is again governed by the loss parameter $p_{\text{loss}}$. In successful entanglement generation events, the link value is updated to “True”, and a time of $T_{\text{success}}$ is added to the cumulative memory time $T$. An additional memory time of $4$ units is added to $T_{\text{round}}$ for each run of established links to account for two qubits held in memory per round of entanglement generation attempts. The total clock time $T_{\text{clock}}$ increases by a fixed increment $T_{\text{clock-step}}$ per round. The protocol repeats until all N1N-1 links have been successfully established, returning both the total memory time $T$ and the total elapsed time $T_{\text{clock}}$.
@@ 3.1 Comparison of Sequential and Parallel Protocols
We now evaluate and compare the respective performances of sequential and parallel asynchronous protocols in a linear quantum network of total length $50$ km and attenuation length $L_{att}=22$ km. Throughout our analysis, we vary both the number of nodes and the memory dephasing time. As discussed in Section 2.3, key metrics for success are the final fidelity of the distributed Bell state and the hashing rate. The results of our analysis are compiled in Figures 1 and 2.
@@
Figure 1: Comparisons between asynchronous distribution protocols, conducted using fidelity and hashing rate metrics. Protocols are compared in networks with attenuation length $L_{att}=22$km and internode distance L=50/NL=50/Nkm, with varying dephasing times $T_{dp}$ and number of nodes NN. Subfigures (a) and (b) give the fidelity of distributed Bell pairs in the sequential and parallel protocols, respectively. Subfigures (c) and (d) depict corresponding hashing rates. While fidelity is comparable across both protocols, the parallel approach demonstrates a strong advantage in hashing rate.
@@
In this section, we presented an analysis comparing the performance of sequential and parallel asynchronous entanglement distribution protocols. We evaluated the output fidelity of distributed Bell pairs, with increasing network size and increasing quantum memory dephasing time, as well as the hashing rates achieved by each protocol. While both approaches produced similar fidelities, the parallel protocol demonstrated a clear advantage in hashing rate due to its overall shorter runtime. In the next section we consider the effect of imperfect Bell state measurement on entanglement distribution, and assess its impact on fidelity and hashing rate.